OpenStudy (anonymous):

Given 3xy-3(x+y)^3=-15 find the equation of the tangent line at the point (2,-3) we have to first use implicit different but...:S

OpenStudy (dumbcow):

this is what i get: \[\frac{dy}{dx} = \frac{y - 3(x+y)^{2}}{3(x+y)^{2} - x}\] at point (2,-3) dy/dx = slope = -6 y + 3 = -6(x -2) y = -6x +9

OpenStudy (anonymous):

using implicit differentiation...

OpenStudy (dumbcow):

yes, then use little algebra to isolate dy/dx

OpenStudy (anonymous):

what'd you get for the implicit differentiation?

OpenStudy (dumbcow):

\[3(\frac{dy}{dx}x + y) - 9(x+y)^{2}(\frac{dy}{dx} +1) = 0\]

OpenStudy (anonymous):

so how'd you get what you got up there?

OpenStudy (dumbcow):

move 2nd term over, divide by 3 on both sides \[\frac{dy}{dx}x + y = 3(x+y)^{2}(\frac{dy}{dx} +1)\] distribute right side \[\frac{dy}{dx}x + y =3\frac{dy}{dx}(x+y)^{2} + 3(x+y)^{2}\] get dy/dx 's on same side \[y - 3(x+y)^{2} = 3\frac{dy}{dx}(x+y)^{2} - \frac{dy}{dx}x\] factor out dy/dx, then divide

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i got it lol! thanks!1

OpenStudy (dumbcow):

:)